Crispy Fried Chicken bought equipment on January 22​, 2018, for $33,000. The equipment was expected to remain in service for four years and to operate for 6,750 hours. At the end of the​ equipment's useful​ life, Crispy estimates that its residual value will be $6,000. The equipment operated for 675 hours the first​ year, 2,025 hours the second​ year,2,700 hours the third​ year, and 1,350 hours the fourth year.

Requirement 1. Prepare a schedule of depreciation​ expense, accumulated​ depreciation, and book value per year for the equipment under the three depreciation​ methods: straight-line,​ units-of-production, and​ double-declining-balance. Show your computations. ​Note: Three depreciation schedules must be prepared.

Begin by preparing a depreciation schedule using the​ straight-line method.

Straight-Line Depreciation Schedule

Depreciation for the Year

Asset

Depreciable

Useful

Depreciation

Accumulated

Book

Date

Cost

Cost

Life

Expense

Depreciation

Value

1-2-2018

12-31-2018

/

4 years

=

12-31-2019

/

4 years

=

12-31-2020

/

4 years

=

12-31-2021

/

4 years

=

27,000

6,000
Answer #1

Ans:

Method 1: Straight Line Method:

The annual depreciation under straight line method is calculated as follows:

Annual Depreciation = (Cost - Salvage Value)/Estimated Life = (33,000 - 6,000)/4 = $6,750

Depreciation Rate=1/life of asset=1/4 years=0.25=25%

The schedule is given below:

Straight-Line Depreciation Schedule

Depreciation for the Year

Asset

Depreciable

Useful

Depreciation

Accumulated

Book

Date

Cost

Cost

Life

Expense

Depreciation

Value

02-Jan-2018

33000

0

0 0

31-Dec-2018

27,000

25%

/

4 years

=

6,750

6,750

26,250

31-Dec-2019

27,000

25%

/

4 years

=

6,750

13500

19,500

31-Dec-2020

27,000

25%

/

4 years

=

6,750

20250

12,750

31-Dec-2021

27,000

25%

/

4 years

=

6,750

27,000

6,000

Book Value=Actual Asset Cost- Accumulated Depreciation